tan2 theta * tan4 theta= 1
find the value of tan3 theta
 

Grade:10

1 Answers

Vikas TU
6 years ago
Dear Student,
tan2θtan4θ = 1.
tan6θ = (tan4θ + tan2θ)/(1 – tan2θtan4θ) -> infinity.
Therefore 6θ = π/2, and 3θ = π/4. Therefore tan3θ = 1.
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)